Two coherent sources of light having intensity ratio 16:1 produce an interference fringe pattern. Calculate the ratio of the intensities of bright and dark fringes.
1
Expert's answer
2020-02-07T10:35:06-0500

Let the intensities of individual source be "16 I\ \ \&\ I"


Bright fringe (maximum) intensity,

"\newline

I_{max} = (\sqrt{I_1 } + \sqrt{I_2})^2"


Dark fringe (minimum) intensity,

"I _\n{min}\n\u200b\t\n =(\\sqrt \nI \n_1\n\u200b\t\n \n\u200b\t\n \u2212 \\sqrt\nI \n_2\n\u200b\t\n \n\u200b\t\n ) \n^2"


"\\dfrac {I_{ \nmax}}\t\n {\nI _{\nmin}\n\u200b}\n\t\n \n\u200b\t\n =\\dfrac{\n( \\sqrt\nI _\n1\t\n + \\sqrt\nI _\n2\t\n ) ^\n2}\n {\n( \\sqrt\nI \n_1\t\n - \\sqrt\nI _\n2\t\n ) ^\n2\n }\n\u200b"

"\implies" "\dfrac{I_{max}}{I_{min}}={\dfrac{(4+1)^2} {(4-1)^2}}=25:9"